I cannot solve this equation: t^2 - 1/6 = 35/6. Thank you for any help.
Answer by mananth(16946)

t^2 - 1/6 = 35/6
t^2=35/6 + 1/6
t^2=36/6
t= sqrt 36/6
t=6/ sqrt 6
t= sqrt6 * sqrt6 / sqrt6
t= sqrt6